If demand and supply are represented by the equations

Demand: LD =13 − (1/3)W

Supply: LS = (1/2) W − 2

then the equilibrium wage (W* ) and employment level (L* ) will equal

W* = $18, L* = 6.

W* = $18, L* = 7.

W* = $16, L* = 8.

W* = $12, L* = 4.
Answer #1

the equilibrium wage (W* ) and employment level (L* ) is where LD =LS

LD = LS

13 -(1/3)W = (1/2)W - 2

15 = 5*W / 6

W* = 15*6 /5

W* = $18

LD = LS at W* =18,

so

L* = (1/2)*18 -2 = 7

L* =7

ANs is b) W* = $18, L* = 7.